APPRENTICE and SIT-INS

Name (33)  @uchicago.edu
William Cole Abram  abramwc
Dylan Allegretti  allegretti
Adam Anderson  adamanderson
Elan Bechor  ebechor
Michael Calderbank  Mcalderbank
Daniel Cervone  dcervone
Beau Dabbs  bdabbs13
Catherine Easton  catherine
Mufti Nomaan Farooqui  mnomy88
Hannah Gordon  hannahg
Austin Gross  Agross
Eric Daniel Jaffe  ericjaffe
Ethan Jerzak  jerzak
Michael (Ryan) Julian  mrjulian
Katayun J. Kamdin  kkamdin
Suyeon Khim  skhim
Satory Kudose  skudose
William Leeb  weleeb
Chenchuan Li  chenchuan
Cong Han Lim  conghan
Ellen McCullagh  ellenm
Scott Messick  smessick
Jamie Morgenstern  jamiemmt
Matt Olson  molson2
Bo Peng  bopeng
Andrew Poppick  apoppick
James Salvatore  jsal13
Honi Sanders  honiraphael
Brad Sugarman  bvs
Michael Travis  mtravis
Brooke Ullery  bullery
Gideon Tan Guang Yuan  serapuff
Tengren Zhang  tengren85

SESAME

Name (22)  @uchicago.edu
Stephan Balady  sbalady
Sarah Bennett  sbennett
David Berger  iloveyou
Chad Casarotto  chadmbol
Christopher Logan Chariker  chariker
Le Chen  lelechen
David Coley  dcoley
Andrew Eisenberg  aeisen
Arpit Gupta  arpit
Adam Kaye  adamkaye
R. Andrew Kreek  rakreek
Joshua Blaise Leners  leners
Carmel Levy  carmel1
Yana Peysakhovich  yana
Robert Prag  rprag
Edward Ruggeri  ruggeri
Maxwell Shron  mshron
Andries Smith  andries
Michael Thvedt  cynicalk
Alexander Volfovsky  alexvolf
Matthew Woolf  mwoolf
Qian Sophia Zhang  qian

COLLEGIATE SCHOLARS

Name (2)  @uchicago.edu
Fan Fei Chong  ffclark
Calvin Lin Zhiwei  linzw

YSP: 7–8

Name (8)  @uchicago.edu
Emily Bargar  ebargar
Padraic James Bartlett  padraic
Sarah Constantin  sconsta1
John Dyer  dyer
Gabriel Gaster  eulerian
Alice Mark  amark
James Marrone  jmar
Brian Taylor  bgtaylor

YSP: 9–10

Name (8)  @uchicago.edu
Daniel Gardner  dbgardner
Drew A. Hudec  dhudec
Nicholas Ford  nford
Grant Larsen  gel
Jean Morrison  jeanm
Alexander J. Munk  amunk
Prerna Nadathur  pnadathur
Caitlin Waddle  kwaddle

YSP: 11–12

Name (6)  @uchicago.edu
Harry Altman  haltman
Corrin Clarkson  clarkson
David Cohen  dcohen1
Lucas Culler  luccul
Jonathan Nieder  jrnieder
Sam Raskin  sraskin